# Privacy-Aware Detection of Fake Identity Documents: Methodology, Benchmark, and Improved Detection Methods (FakeIDet2)

### v1

@@ 4.1 Patch Embedding Extractor
To learn features at patch level effectively, we go deep into representation learning [6, 25, 9], aiming to obtain compressed, meaningful representations from high-dimensional data into the embedding space. To do so, we treat patches coming from both real and fake IDs as four different classes (i.e., 3 PAI s + real) and use DINOv2’s (ViT-S/14) [26] backbone with its weights frozen. Using DINOv2 for extracting patch embeddings seems appropriate, since the goal for its original training procedure was to match the embedding distribution between images where the full context is given and random patches extracted from the same image. To leverage these features, we attach a final layer to DINOv2 embeddings, hence, our Patch Embedding Extractor is defined as an encoder function that maps an input patch $x_{i}\in\mathbb{R}^{p\times p\times c}$, where $p$ is the patch size and $c$ is the channel dimension (i.e., $c=3$ in the RGB spectrum), to its corresponding compressed representation $z_{i}\in\mathbb{R}^{d}$, where $d$ is the dimension of the embedded space.
We explore several softmax-based loss functions, i.e., CosFace [38], ArcFace [10], and AdaFace [17], to asses which one suited our problem best for discriminative representation learning. The generic formulation for the Softmax loss function is defined in Eq. 1, where $C$ is the number of classes and $f(\theta_{y_{i}})$ is a function that computes the alignment between the embedding obtained from the input image and the class prototype of its label $y_{i}$. Hence, the objective for an embedding of one class is to be as close as possible to other embeddings from the same class (better aligned) while being apart from embeddings of other classes by an imposed distance, called margin (worst aligned). The main differences between these loss functions reside in how the margin is computed for the alignment function, $f(\theta_{y_{i}})$. For example, CosFace, defined in Eq. 2, computes the loss by adding the margin $m$ to the cosine similarity, while ArcFace, defined in Eq. 3, computes the loss by adding an angular margin $m$ to the cosine similarity. In the case of AdaFace, defined in Eq. 4, the margin is computed adaptively, depending on the image quality, which is computed as the norm of embedding obtained from the input image, which is normalized based on batch statistics for better stability, which is denoted by $|\widehat{z_{i}}$$|$.
$\mathcal{L}=\log\left(\frac{\text{exp}(f(\theta_{y_{i}}))}{\text{exp}(f(\theta_{y_{i}}))+\sum_{j\neq y_{i}}^{C}\text{exp}(s\cos\theta_{j})}\right)$ (1)
@@
One issue that we encountered when training the Patch Embedding Extractor was that softmax-based loss functions did not account for the severe class imbalance in the physical composite attacks (see Table 3). To address this limitation, we propose to apply dynamic class weights to these loss functions, which evolve as the training progresses. The initial weights $w_{0}\in\mathbb{R}^{C}$, where $C$ is the number of classes, are computed as the normalized inverse of the frequency for each class using Eq. 5, where $N$ is the total number of patches in the database, and $N_{i}$ is the total number of patches belonging to the i-th class.
@@
After that, the class weights are updated at each epoch until they all converge to 1, hence, giving progressively the same weight to all classes. The update rule is presented in Eq. 6, where $w_{t}\in\mathbb{R}^{C}$ is a vector that stores the class weights at time-step $t$, $\mathbf{1}\in\mathbb{R}^{C}$ is a vector full of ones, $\lambda_{t}\in[0,1]$ is a pondering factor that linearly decays from 1 to 0 as training progresses, and $e$ is the total number of time-steps involved in the training procedure. Intuitively, at the beginning of the training, the class weights are going to be close to $w_{0}$ giving more importance to the under-represented class, but, as time-steps go by, the class weights are going to get closer to 1, giving all classes the same importance by the end of the training.
@@
These class-weights are then used to modulate the gradient signal that the margin loss function produces, giving higher values when the embeddings from the underrepresented class are close to the embeddings from another class. The modified version of this loss is defined in Eq. 7, where $\mathcal{L}$ is the loss function defined in Eq. 1, $w_{t,y_{i}}$ is the class weight associated to the label $y$ of the $i$-th element of the batch at epoch $t$.
@@ 4.2 Patch Embedding Fusion
To effectively fuse all patch embeddings of a given ID, we make use of the MHSA layer presented in the Transformer architecture [36]. This mechanism was introduced to solve issues related to model long-range dependencies, which are typical in Recurrent Neural Neworks (RNN) models, which process information sequentially. However, in the Self-Attention mechanism, all the elements of the input sequence are processed at the same time, enabling modeling dependencies from tokens that are from afar. The Self-Attention layer expects as input a sequence of embeddings (or tokens) $X\in\mathbb{R}^{n\times d}$, where $n$ is denoted as the sequence length and $d$ the embedding dimension. The sequence is processed according to the Self-Attention mechanism, defined in Eq. 8, where $Q$ is the query, $K$ is the key, $V$ is the value, and $d_{k}$ is the projected embedding dimension for the query and the key. The query, key and value matrices are single linear projections from $X$, which are computed as $Q=XW_{Q}$, $K=XW_{K}$ and $V=XW_{V}$, respectively, where $W_{Q}\in\mathbb{R}^{d\times d_{k}}$, $W_{K}\in\mathbb{R}^{d\times d_{k}}$ and $W_{V}\in\mathbb{R}^{d\times d_{v}}$, where $d_{v}$ is the projected dimension for the value. Hence, the obtained sequence after the Self-Attention mechanism is a matrix $X^{\prime}\in\mathbb{R}^{n\times d}$ where each token values have been modified depending on their importance in the sequence, compared to their peers.
@@
Furthermore, the attention mechanism allows the queries, keys, and values to be linearly projected $h$ times using different learned projections of dimensions $d_{k}$ and $d_{v}$. To achieve this, $d_{k}$ and $d_{v}$ are typically set to $d/h$, where $d$ is the model dimension. Each of these computations defines an attention head (Eq. 9). The outputs of the $h$ attention heads are concatenated and projected back to the original input dimension using a final linear transformation $W^{O}\in\mathbb{R}^{hd_{v}\times d}$ (Eq. 10).
@@
Since the Patch Embedding Extractor produces an embedding $z_{i}$ for each patch $x_{i}$, we treat the set of patches that come from a single ID as a sequence $X=[z_{0};z_{1};...;z_{n}]\in\mathbb{R}^{n\times d}$, where $n$ is the number of patches from a single ID and $d$ is the dimension from the embedding obtained from the Patch Embedding Extractor. These embeddings are then normalized via layer normalization [3] and forwarded into the MHSA layer, that produces a feature sequence $X^{\prime}=[z_{0}^{\prime};z_{1}^{\prime};...;z_{n}^{\prime}]\in\mathbb{R}^{n\times d}$ where the embeddings values $z_{i}^{\prime}$ have been modified according to their importance in the input sequence. Additionally, we leverage residual connections [14] to enable better gradient flow in the training procedure, as two layers of MHSA and layer normalization are stacked together in this module, where the first MHSA layer has $h=8$ attention heads and the second layer has $h=4$ attention heads.
@@ 5 Experimental Setup
In order to define a standard, reproducible benchmark that allows to advance properly in this research field, in this section we define our experimental setup. First, Sec. 5.1 describes the specific hyperparameters of our proposed FakeIDet2 method. Then, in Sec. 5.2 we describe the experimental protocol and the different scenarios of interest in the analysis. Finally, Sec. 5.3 describes the evaluation metrics. As a general note, we use the EER metric as a performance indicator to obtain the best patch and anonymization configuration. For all training procedures and experiments, we used only one NVIDIA RTX 3080 with 10GB of VRAM.
@@ 5.1 FakeIDet2: Hyperparameters
Regarding the Patch Embedding Extractor, for any patch size the input is resized to $224\times 224$, which is the input shape of DINOv2, with an embedding layer of dimension $d=128$. We use Adam optimizer with weight decay, an initial learning rate $\alpha_{0}=0.00125$, and exponential decays for the momentum estimators of $\beta_{1}=0.9$ and $\beta_{2}=0.999$, respectively, and a weight decay $\delta=0.0001$. All models are trained for 70 epochs, using the modified versions of the softmax-based losses presented in Sec. 4.1 with an early stopping condition that ends the training if the loss in the validation does not decrease for 5 consecutive epochs. Additionally, we use a cosine annealing decay scheduler for the learning rate, with values decreasing from $\alpha_{0}=1.25e-3$ to $\alpha_{e}=1.25e-4$, and a batch size of 256. Regarding data transformations, we opt to randomly apply Gaussian blur with a kernel size $k=3$ and color jitter with a brightness of 0.2, a hue of 0.05, a contrast of 0.25, and a saturation of 0.2. These data transformations are applied in training time with a probability of $p=0.2$. Finally, the pixel values are normalized according to ImageNet mean and standard deviation statistics.
For the Patch Embedding Fusion, in particular the MHSA layers, we define a context length of 384 for both patch configurations, which is enough so that no patch from an ID is left out when performing a forward pass. Furthermore, we use masked-MHSA to indicate which embeddings should be attended, so that in case the number of patches from a single ID is fewer than the context length, the introduced padded embeddings are ignored. We use Binary Cross Entropy (BCE) loss with class weights, since there is a class ratio of 3 to 1 between fake and real data. For all of our experiments, we choose Adam as the optimizer, using an initial learning rate $\alpha_{0}=1.25e-4$ and exponential decays for the momentum estimators of $\beta_{1}=0.9$ and $\beta_{2}=0.999$, respectively, with 1 epoch of learning rate warm-up. Additionally, the learning rate followed a cosine annealing decay from $\alpha_{0}=1.25e-4$ to $\alpha_{e}=1.25e-5$.
@@ 5.2 Experimental Protocol
In order to assess the feasibility of our proposed FakeIDet2 detector and privacy-aware scenarios to facilitate the collaboration between ID Holders and AI Resesearchers, we propose a novel experimental protocol that aims to: i) evaluate FakeIDet2 performance regarding patch size and anonymization configurations in our novel FakeIDet2-db, comparing the results with the traditional case in the literature, i.e., feeding the fake detectors with the whole non-anonymized ID, and ii) explore domain adaptation, by evaluating FakeIDet2 detector on popular databases in the topic of fake ID detection.
@@
First, in all experiments that involve any kind of training procedure, the database is divided into a development set (80% of the real and fake IDs) and a final evaluation set (the remaining 20%) for all training scenarios. This split simulates a realistic scenario in which the evaluation involves unseen identities. In addition, for our FakeIDet2-db, the first version of the Spanish ID is included only in the final evaluation set, not in the development set, in order to analyze the generalization ability of FakeIDet2 to other ID templates.
@@
Regarding the experiments included in Sec. 6, we first validate the technical details of our proposed FakeIDet2 in terms of performance and privacy. Sec. 6.1 explores the different softmax-based loss functions for representation learning, which will determine the best loss function for the Patch Embedding Extractor module. After that, an exploration in terms of the optimal patch size is conducted in Sec. 6.2, to test whether more patches at lower sizes (i.e., $64\times 64$) yield better performance compared to fewer patches of greater size (i.e., $128\times 128$). Additionally, we also compare the performance of our proposed FakeIDet2 in terms of anonymization levels in Sec. 6.3, which also goes deep into interesting scenarios such as how well a model trained on fully-anonymized data performs on non-anonymized data. Finally, for completeness, we compare FakeIDet2 with our previous FakeIDet [23]. From this set of experiments we will select the best configuration from FakeIDet2 in terms of privacy versus performance.
@@
Then, after validating the technical contributions of FakeIDet2, we analyze its generalization capabilities against unseen scenarios, including leave-one-out experiments for two interesting scenarios: novel attack detection and out-of-distribution sensors. In this case, we train FakeIDet2 with the optimal configuration, leaving either images belonging to one PAI or images taken from one sensor out of the training data, testing the performance of FakeIDet2 when exposed to unseen attacks, or images taken from acquisition devices with sensors that were not used during training. The analysis of these scenarios are available in Sec. 6.4 and Sec. 6.5, respectively.
@@
Finally, we evaluate in Sec. 6.6 the best configuration of FakeIDet2 in a cross-database scenario, providing a standard public benchmark to the research community. This scenario intends to evaluate our proposed model using unseen databases, where IDs are from different countries, including different templates, and are acquired under different conditions in terms of illumination and acquisition devices. Also, these databases contain different PAIs compared to our proposed database. With this goal in mind, we select three different databases for performance evaluation, DLC-2021 [30], KID34K [27] and Benalcazar et al. [5]. These databases are selected for the following reasons: i) each database has ID images from three different regions of the world (i.e., Europe, Asia and South America), hence the evaluation is conducted on different distributions in terms of ID templates with different alphabets, and ii) we contemplate physical attacks as well as synthetic attacks in our evaluation, to assess the performance of FakeIDet2 to synthetic data in the form of fake IDs.

### v2

@@ 4.1 Patch Embedding Extractor
To effectively learn features at patch level, we go deep into representation learning, aiming to obtain compressed, meaningful representations from high-dimensional data into the embedding space. To do so, we treat patches coming from both real and fake IDs as four different classes (i.e., 3 PAI s + real) and use DINOv2’s (ViT-S/14) [27] backbone with its weights frozen. Using DINOv2 for extracting patch embeddings seems appropriate, since the goal of its original training procedure was to match the embedding distribution between images where the full context is given and random patches extracted from the same image. To take advantage of these features, we attach a final layer to DINOv2 embeddings; hence, our Patch Embedding Extractor is defined as an encoder function that maps an input patch $x_{i}\in\mathbb{R}^{p\times p\times c}$, where $p$ is the patch size and $c$ is the channel dimension (i.e., $c=3$ in the RGB spectrum), to its corresponding compressed representation $z_{i}\in\mathbb{R}^{d}$, where $d$ is the dimension of the embedded space.
We explore three softmax-based loss functions [23]: CosFace [38], ArcFace [11], and AdaFace [17], to assess which one best fits our problem to learn discriminative representations. The generic formulation for the softmax loss function is defined in Eq. 1, where $C$ is the number of classes, $f(\theta_{y_{i}})$ is a function that computes the alignment between the embedding obtained from the input image and the class prototype of its label $y_{i}$ and $s$ is a scaling factor that modulates the alignment value. Hence, the objective for an embedding of one class is to be as close as possible to other embeddings from the same class (better aligned) while being apart from embeddings of other classes by an imposed distance, called margin (worst aligned). The main differences between these loss functions reside in the way the margin is computed for the alignment function, $f(\theta_{y_{i}})$. For example, CosFace, defined in Eq. 2, computes the loss by adding the margin $m$ to the cosine similarity, while ArcFace, defined in Eq. 3, computes the loss by adding an angular margin $m$ to the cosine similarity. In the case of AdaFace, defined in Eq. 4, the margin is adaptively calculated, depending on the quality of the image [31], which is calculated as the embedding norm obtained from the input image normalized on batch statistics for better stability and is denoted by $|\widehat{z_{i}}$$|$.
$\mathcal{L}=\log\left(\frac{\text{exp}(f(\theta_{y_{i}}))}{\text{exp}(f(\theta_{y_{i}}))+\sum_{j\neq y_{i}}^{n}\text{exp}(s\cos\theta_{j})}\right)$ (1)
@@
An issue we encountered when training the Patch Embedding Extractor was that softmax-based loss functions did not account for the severe class imbalance in the physical composite attacks (see Table 3). To address this limitation, we propose to apply dynamic class weights to these loss functions that evolve as training progresses [14]. The initial weights $w_{0}\in\mathbb{R}^{C}$, where $C$ is the number of classes, are calculated as the normalized inverse of the frequency for each class using Eq. 5, where $N$ is the total number of patches in the database and $N_{i}$ is the total number of patches belonging to the $i$-th class.
@@
After that, the class weights are updated at each epoch until they all converge to 1, hence progressively giving the same weight to all classes. The update rule is presented in Eq. 6, where $w_{t}\in\mathbb{R}^{C}$ is a vector that stores the weights of the class at the time step $t$, $\mathbf{1}\in\mathbb{R}^{C}$ is a vector full of ones, $\lambda_{t}\in[0,1]$ is a pondering factor that linearly decays from 1 to 0 as training progresses, and $e$ is the total number of time steps involved in the training procedure and $\lambda_{e}=0$. Intuitively, at the beginning of the training, the class weights will be close to $w_{0}$, giving more importance to the underrepresented class, but, as time goes by, the class weights will be closer to 1, giving all classes the same importance at the end of the training.
@@
These class weights are then used to modulate the gradient signal that the margin loss function produces, giving higher values when the embeddings of the underrepresented class are close to the embeddings from another class. The modified version of this loss is defined in Eq. 7, where $\mathcal{L}$ is the loss function defined in Eq. 1, $w_{t,y_{i}}$ is the class weight associated with the label $y$ of the $i$-th element of the batch at the time $t$.
@@ 4.2 Patch Embedding Fusion
To effectively fuse all patch embeddings of a given ID, we use a Multi-Head Self-Attention (MHSA) layer [10]. This self-attention mechanism was introduced to solve problems related to long-range dependency issues of the model, which are typical in Recurrent Neural Neworks (RNN) models, which process information sequentially. However, in the self-attention mechanism, all elements of the input sequence are processed at the same time, enabling modeling of dependencies from distant tokens. The Self-Attention layer expects as input a sequence of embeddings (or tokens) $X\in\mathbb{R}^{n\times d}$, where $n$ is denoted as the sequence length and $d$ the embedding dimension. The sequence is processed according to the Self-Attention mechanism, defined in Eq. 8, where $Q$ is the query, $K$ is the key, $V$ is the value, and $d_{k}$ is the projected embedding dimension for the query and the key. The query, key, and value matrices are single linear projections of $X$, which are calculated as $Q=XW_{Q}$, $K=XW_{K}$ and $V=XW_{V}$, respectively, where $W_{Q}\in\mathbb{R}^{d\times d_{k}}$, $W_{K}\in\mathbb{R}^{d\times d_{k}}$ and $W_{V}\in\mathbb{R}^{d\times d_{v}}$, where $d_{v}$ is the projected dimension of the value. Hence, the sequence obtained after the self-attention mechanism is a matrix $X^{\prime}\in\mathbb{R}^{n\times d}$ where each token value has been modified depending on their importance in the sequence, compared to their peers.
@@
Furthermore, the attention mechanism allows queries, keys, and values to be linearly projected $h$ times using different learned projections of dimensions $d_{k}$ and $d_{v}$. To achieve this, $d_{k}$ and $d_{v}$ are typically set to $d/h$, where $d$ is the dimension of the model. Each of these computations defines an attention head (Eq. 9). The outputs of the $h$ attention heads are concatenated and projected back to the original input dimension using a final linear transformation $W^{O}\in\mathbb{R}^{hd_{v}\times d}$ (Eq. 10).
@@
Since the Patch Embedding Extractor produces an embedding $z_{i}$ for each patch $x_{i}$, we treat the set of embeddings that come from a single ID as a sequence $Z=[z_{1};z_{2};...;z_{n}]\in\mathbb{R}^{n\times d}$, where $n$ is the number of patches from a single ID and $d$ is the dimension of the embedding obtained from the Patch Embedding Extractor. These embeddings are then normalized through layer normalization [3] and are transmitted to the MHSA layer that produces a feature sequence $Z^{\prime}=[z_{1}^{\prime};z_{2}^{\prime};...;z_{n}^{\prime}]\in\mathbb{R}^{n\times d}$ where the embedding values $z_{i}^{\prime}$ have been modified according to their importance in the input sequence. Furthermore, we leverage residual connections to enable a better gradient flow in the training procedure, since two layers of MHSA and layer normalization are stacked in this module, where the first MHSA layer has $h=8$ attention heads and the second layer has $h=4$ attention heads.
@@ 5 Experimental Setup
In order to define a standard, reproducible benchmark that allows us to advance properly in this research field, in this section, we define our experimental setup. First, Sec. 5.1 describes the specific hyperparameters of our proposed FakeIDet2 method. Then, in Sec. 5.2 we describe the experimental protocol and the different scenarios of interest in the analysis. Finally, Sec. 5.3 describes the evaluation metrics. In general, we use the EER metric as a performance indicator to obtain the best patch and anonymization configuration. For all training procedures and experiments, we used only one NVIDIA RTX 3080 with 10GB of VRAM.
@@ 5.1 FakeIDet2: Hyperparameters
Regarding the Patch Embedding Extractor, for any patch size, the input is resized to $224\times 224$, which is the input shape of DINOv2, with an embedding layer of dimension $d=128$. We use Adam optimizer with weight decay, an initial learning rate $\alpha_{0}=0.00125$, and exponential decays for the momentum estimators of $\beta_{1}=0.9$ and $\beta_{2}=0.999$, respectively, and a weight decay $\delta=0.0001$. The modified versions of the softmax-based losses presented in Sec. 4.1 are defined with a margin of $m=0.4$, which are the optimal experimental values that perform well for low-quality datasets without compromising the capacity in high-quality data, and a scaling factor of $s=64$, as it provides a strong signal from both samples that are close and far from the decision boundary, creating better separability in the learned representations according to [17]. All models are trained for 70 epochs, with an early stopping condition that ends the training if the loss in the validation does not decrease for 5 consecutive epochs. Furthermore, we use a cosine annealing decay scheduler for the learning rate, with values decreasing from $\alpha_{0}=1.25e-3$ to $\alpha_{e}=1.25e-4$, and a batch size of 256. Regarding data transformations, we opt to apply random Gaussian blur with a kernel size $k=3$ and color jitter with a brightness of 0.2, a hue of 0.05, a contrast of 0.25, and a saturation of 0.2. These data transformations are applied in training time with a probability of $p=0.2$. Finally, the pixel values are normalized according to ImageNet mean and standard deviation statistics.
For the Patch Embedding Fusion, in particular the MHSA layers, we define a context length of 384 for both patch configurations, which is enough so that no patch from an ID is left out when performing a forward pass. Furthermore, we use masked-MHSA to indicate which embeddings should be attended, so that in case the number of patches from a single ID is fewer than the context length, the introduced padded embeddings are ignored. We use Binary Cross Entropy (BCE) loss with class weights, since there is a class ratio of 3 to 1 between the fake and the real data. For all of our experiments, we choose Adam as optimizer, using an initial learning rate $\alpha_{0}=1.25e-4$ and exponential decays for the momentum estimators of $\beta_{1}=0.9$ and $\beta_{2}=0.999$, respectively, with 1 epoch of learning rate warm-up. Furthermore, the learning rate followed a cosine annealing decay from $\alpha_{0}=1.25e-4$ to $\alpha_{e}=1.25e-5$.
@@ 5.2 Experimental Protocol
In order to assess the feasibility of our proposed FakeIDet2 detector and privacy-aware scenarios to facilitate collaboration between ID Holders and AI Researchers, we propose an experimental protocol that aims to: i) evaluate FakeIDet2 performance regarding patch size and anonymization configurations in our new FakeIDet2-db, comparing the results with the traditional case in the literature, i.e., feeding the fake detectors with the whole non-anonymized ID, and ii) explore domain adaptation, by evaluating FakeIDet2 on popular databases in the topic of fake ID detection.
@@
First, in all experiments that involve any kind of training procedure, the database is divided into a development set (80% of the real and fake IDs) and a final evaluation set (the remaining 20%) for all training scenarios. This split simulates a realistic scenario in which the evaluation involves unseen identities. In addition, for our FakeIDet2-db, the first version of Spanish ID is included only in the final evaluation set, not in the development set, to analyze the generalizability of FakeIDet2 to other ID templates.
@@
Regarding the experiments included in Sec. 6, we first validate the technical details of our proposed FakeIDet2 in terms of performance and privacy. Sec. 6.1 explores the different softmax-based loss functions for representation learning, which will determine the best loss function for the Patch Embedding Extractor module. After that, an exploration in terms of the optimal patch size is conducted in Sec. 6.2, to test whether more patches at lower sizes (i.e., $64\times 64$) yield better performance compared to fewer patches of greater size (i.e., $128\times 128$). Additionally, we also compare the performance of our proposed FakeIDet2 in terms of anonymization levels in Sec. 6.3, which also goes deep into interesting scenarios such as how well a model trained on fully-anonymized data performs on non-anonymized data. Finally, for completeness, we compare FakeIDet2 with our previous FakeIDet [25]. From this set of experiments, we will select the best configuration of FakeIDet2 in terms of privacy versus performance.
@@
Then, after validating the technical contributions of FakeIDet2, we analyze its generalization capabilities against unseen scenarios, including leave-one-out experiments for two interesting scenarios: novel attack detection and out-of-distribution sensors. In this case, we train FakeIDet2 with the optimal configuration, leaving images belonging to one PAI or images taken from one sensor out of the training data, testing the performance of FakeIDet2 when exposed to unseen attacks, or images taken from acquisition devices with sensors that were not used during training. The analyses of these scenarios are available in Sec. 6.4 and Sec. 6.5, respectively.
@@
Finally, we evaluate in Sec. 6.6 the best configuration of FakeIDet2 in a cross-database scenario, providing a standard public benchmark for the research community. This scenario aims to evaluate our proposed model using unseen databases, where the IDs are from different countries, including different templates, and are acquired under different conditions in terms of illumination and acquisition devices. In addition, these databases contain different PAIs compared to our proposed database. With this goal in mind, we select three different databases for performance evaluation, DLC-2021 [30], KID34K [28] and Benalcazar et al. [5]. These databases are selected for the following reasons: i) each database has ID images from three different regions of the world (i.e., Europe, Asia and South America); hence the evaluation is conducted on different distributions in terms of ID templates with different alphabets, and ii) we contemplate physical attacks as well as synthetic attacks in our evaluation, to assess the performance of FakeIDet2 to synthetic data in the form of fake IDs.